OpenStudy (anonymous):

Calculate the Riemann sum R(f,P,C) for f(x)=x, P= {1,1.3,1.7,2}, C={1.2,1.5,2}

OpenStudy (anonymous):

i guess P is the partition what is C?

OpenStudy (anonymous):

oh i bet C is the points at which to evaluate the function right?

OpenStudy (anonymous):

I'm not sure, all that was given was what I copied and pasted.

OpenStudy (anonymous):

not to be a jerk, but if you don't understand that the notation means, then you cannot start the problem i have not seen this notation before, but i imagine that your job is to evaluate the function at the points in C and multiply each by the length of each interval

OpenStudy (anonymous):

from 1 to 1.3 is length .3, and \(f(1.2)=1.2\) so the first term in the sum will be \[1.2\times .3\]

OpenStudy (anonymous):

then from 1.3 to 1.7 is length .4, and \(f(1.5)=1.5\) so the second term in the sum will be \[1.5\times .4\]

OpenStudy (anonymous):

I'm sorry that I don't understand the question? That was why I was asking on here, obviously. Thanks for your help but I don't need disrespectful answers, because regardless of if you preface something with "not to be a jerk/not to be rude" that doesn't change the fact that the statement was rude. Thank you for your help I just don't appreciate being talked down to.

OpenStudy (anonymous):

and finally from 1.7 to 2 is length .3 and \(f(2)=2\) so the last term was \(2\times .3\)

OpenStudy (anonymous):

Thank you again for your help.